# Hertzsprung gap stars in nearby galaxies and the Quest for Luminous Red Novae Progenitorsthanks: Tables 3 and 4 are only available in electronic form at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/cgi-bin/qcat?J/A+A/

### v1

@@ 3.2 Contamination fraction
To evaluate the contamination fraction from intrinsic extinction and foreground sources, we considered various types of contaminants. Foreground contamination was quantified using the TRILEGAL star counts. According to the method detailed in Section 2.4, approximately 5% of the sources in our YSG sample are expected to be foreground contaminants (including sources having high Gaia proper motion). The foreground contamination mainly affects galaxies at low Galactic latitudes ($|b|<25$º) and 75% of the galaxies have less than 10% of their YSG candidates being foreground, as shown in Figure 10 (right panel). As illustrated in this Figure, galaxies closer to the Galactic plane or toward the Galactic center exhibit higher contamination fractions. After using Gaia proper motions to clean the sample, removing 8,375 sources (3.4% of the sample), we therefore estimate the remaining foreground contamination to be $\sim$ 1.7%.
@@
Finally, contaminants such as quasars, Cepheids and luminous blue variables were identified and removed using Gaia and Simbad databases. These known contaminants represent a minority of the sample across all distances (Figure 7). These sources are typically brighter in both observed and absolute magnitudes, as shown in Figure 12. The clean sample of YSG contains 233,336 sources.
@@ 4.1 Completeness for known LRN progenitors
Table 5: LRN progenitors identified in the literature, and their retrieval status in our YSG sample. References: [1] MacLeod et al. 2017, [2] Pastorello et al. 2021a, [3] Pastorello et al. 2021b, [4] Blagorodnova et al. 2017, [5] Blagorodnova et al. 2021, [6] Cai et al. 2022 [7] Pastorello et al. 2023, [8] Smith et al. 2016a.
Many extragalactic LRNe discovered during the last decade had pre-outburst images taken several years before the transient (e.g. MacLeod et al. 2022 and references therein). For eight of them, listed in Table 5, their host galaxy is in our sample, allowing us to assess the rate of retrieved progenitors in our sample of YSG candidates. Three out of eight progenitors are missing in our sample. For AT2015dl, the transient is located outside the footprint of the HST observations. For NGC4490-OT, although the progenitor was measured and studied by Smith et al. (2016b), its apparent magnitude of $m_{\mathrm{F606W}}=23.58\pm$0.24 was too faint to be retrieved as an HSCv3 source (the faintest retrieved sources in a 30 arcsec radius circle around the progenitor position were 23.2 mag). For AT2019zhd, the progenitor had a low luminosity ($M_{\mathrm{F555W}}=0.21\pm 0.14$, Pastorello et al. 2021a), placing it in the most populated region of the Hertzsprung gap. Consequently, it could not be selected after the Gaussian mixture cut. Moreover, its proximity to the edge of the field of view prevents this source from being retrieved as a MAST-catalogued source. Considering these results, our sample appears to represent a relatively comprehensive selection of LRN progenitors, with a $\sim$70completeness level in regions covered by HST.
@@ 4.2 Progenitors of past and ongoing transients: crossmatch to TNS
M 31 19.8 2016ept 0.38 Clear 18.7 -5.7
M 31 21.3 2016fbx 0.26 Clear 17.7 -6.7 Nova
NGC 3938 24.2 2017ein 0.56 Clear 17.6 -13.5 SN Ic
@@
Table 6: YSG candidates matching TNS objects. The last column shows the type according to TNS with eventual corrections (V*: variable star, f.a.: false alarm). Objects are sorted by TNS name.
Our YSG sample can be used to find possible progenitors for past transients and further analyse them. To this end, we crossmatched our YSG candidates to the TNS public objects as of 2024 August 31st, with a match radius of 0.6”. The resulting list of 32 transients is detailed in Table 6. About half of these transients are already classified and some are well-studied in the literature (e.g. SN2016bau: Aryan et al. 2021. SN2017ein: Van Dyk et al. 2018; Teffs et al. 2021. SN2024ggi: Pessi et al. 2024; Jacobson-Galán et al. 2024). For those, our HST matches may provide useful pre-outburst luminosity levels. The TNS report of AT2019ejn mentions multiple discoveries of this transient, suggesting it may be an LBV. For three other transients (AT2020aaqy, AT2021tmm and AT2021ytf), the $F814W$ magnitude is brighter than the TNS discovery magnitude, making them likely to also be variable stars. For twelve other transients, the type was not straightforwardly identified in TNS and they deserve a detailed analysis of their light curves and HST progenitors. Therefore, we inspected HST colour images for every candidate, using the Hubble Legacy Archive (HLA) website [^11] https://hla.stsci.edu/hlaview.html. Cutouts of 10 arcsec side-to-side are displayed in Figure 13 and the position of the transient is marked by a circle.
@@
Figure 14: ZTF forced photometry light curves of the 12 candidates resulting from our TNS cross-match. The blue areas indicate the period that was used to set the baseline flux at the location of these transients.
AT2016ept: this sources matches a YSG in M31 MAST catalogues, with $m_{F814W}$ = 19.8, A_F475W = 21.9. After inspection of the HLA image we found that this source is the confusion of several fainter sources ($m_{F814W}>20.4$) that are well listed in HSCv3, although none of them is a YSG. The dips in its light curve could indicate the presence of an eclipsing binary system within this stellar environment, but there is no evidence of LRN precursor emission.
@@
AT2021blt: this transient is located at 0.59” of a $\sim 11M_{\odot}$ YSG candidate, in the middle of the Hertzsprung gap in the CMD. With only 3 detections, no long-term evolution can be detected in its forced photometry light curve. Given the faint peak absolute magnitude, this transient may be a stellar outburst.
AT2021mjv: this transient matches a $\sim 11M_{\odot}$ YSG with a separation of 0.21”. This HST source is located in the Hertzsprung gap, towards the MS in the CMD. It is in a crowded area inside a stellar cluster. Source confusion is highly likely for this candidate, affecting its photometry (CI = 1.8). Given its apparent colour and starforming environment, we argue that this HST source is likely to be a MS star ($P_{\mathrm{extin}}=0.84$) selected as YSG because of source confusion. It may be a long-period variable as perhaps suggested by its light curve.
AT2022llt: this transient matches a $>20M_{\odot}$ YSG with a separation of 0.06”. Although ZTF forced photometry does not reveal any long-term behavior, the ATLAS light curve shows a slow brightening of the source in recent years (Figure 16, top-middle panel). Image inspection shows that source confusion is unlikely for this source (also supported by its CI=1.05). Followup of this source is ongoing and encouraged.
@@ 4.3 Other precursor candidates
Table 7: Precursor candidates (sorted by galaxy name) identified in ZTF, MeerLICHT, or BlackGEM. The last column gives the slope of the best-fit linear trend to the ATLAS light curve.
@@ 5.1 Comparison of our YSG candidates to spectroscopic samples
In addition to using proper motions, variability is an independent criterion that helps distinguishing a YSG from a foreground star. For example, massive YSGs have been known to vary erratically compared to low-mass dwarfs. The bottom panel of Figure 17 shows the distribution of the Gaia-measured dGmag, the difference between brightest and faintest $G$-band magnitude observed during the course of the survey. Interestingly, stars classified as foreground but with a low PM exhibit variability on average 6 times greater than those with proper motion data (average dGmag=0.2 instead of 0.033, Figure 17). This suggests a misclassification of these stars, which may be actual members of M31. To test this hypothesis, we model the distribution of dGmag as a sum of two distributions: one for high-PM sources identified as foreground, and one for low-PM sources identified as YSGs and RSGs. We can then estimate the contribution of YSGs and RSGs within low-PM foreground-classified stars. The fraction of these candidate members is 68 $\pm$ 11%. This suggests that the residual contamination within low-PM stars with $V<18.5$ is actually $\sim$ 8% and not 23%. Furthermore, if our sample were to be used to identify progenitors of variable or transient events, the expected value of dGmag would be relatively large, further reducing the probability of association with a foreground source.
@@
Finally, we can quantify the completeness of our YSG catalogue in different magnitude bins. Using for reference the samples of YSG candidates in M31 from Drout et al. (2009) and Gordon et al. (2016) (see the corresponding CMD in Figure 18), and applying the same proper motion cut as done in our study, we obtain the completeness estimates shown in Figure 19. While the brightest stars are not always detected in third-party HST catalogues (and surveys other than HST are well-adapted to recover those), the fraction of recovered YSGs among recovered HST sources consistently remains around 80%. The lost fraction is primarily located in the dotted-line rectangle shown in Figure 18. This loss occurs because our conservative threshold, designed to exclude the main-sequence and some extinguished main-sequence stars, also omits YSGs in the bluer part of the CMD.
@@ 5.2 Impact of photometric uncertainties and source confusion
One method to quantify source confusion is by using the concentration index (CI), defined as the difference in magnitude between two apertures, typically a smaller aperture of 0.05” and a larger aperture of 0.15” (for ACS and WFC3/UVIS), normalised such that its distribution peaks at 1. Objects with CI values around 1.0 are likely to be stars, while those with significantly higher CI values are likely to be extended sources or the confusion of several stars. Figure 20 presents the distribution of YSG candidates in the CI – $m_{F814W}$ plane. At bright magnitudes, the CI is close to 1, showing that sources are point-like and source confusion is not significant. The faint sources are more numerous and thus denser on the sky, especially in distant galaxies. Therefore, they are more affected by source confusion and can have CI values close to 2.
Similarly, photometric uncertainties can bias the selection of YSGs in HST observations. These uncertainties arise from various sources, such as photon noise, background subtraction errors, and instrumental effects. When dealing with YSG candidates, identified based on their position in the CMD, even small errors in photometry can shift stars into or out of the YSG region. For example, in the NGC 4455 galaxy, an edge-on spiral galaxy located 7.3 Mpc away, observations were conducted using the $F814W$ filter for 1030 seconds. These parameters are close to the median values of our sample of galaxies with YSG candidates. When the colour-magnitude diagram is regenerated using magnitude values and errors modelled as a normal distribution, 9% of YSG candidates shift out of the selection region. Conversely, a similar number of previously unselected stars are shifted into the YSG selection region.
@@ 5.3 Search for LRN Progenitors and Precursors with LSST
Conversely, the study projects that LSST, over its 10-year survey, could detect around 50 red supergiant progenitors and several yellow and blue supergiants. It estimates that progenitors of Type Ic supernovae would be detectable if they are brighter than -4.0 magnitudes in the LSST $i$ band, respectively. Given their similar spectral type (A5) compared to YSGs, we assume a similar performance for YSG detection: we therefore expect that 77% of our YSG sample will be redetected by LSST, for sources having $M_{F814W}<-4$ (or $M_{*}>6M_{\odot}$, Figure 9). This will provide exquisite variability constraints for sources of $r<23$, allowing us to identify new LRN precursors, given LSST repeated observations and the long timescales of these events. Furthermore, LSST can expand the progenitor sample for the 1235 galaxies within 20 Mpc and at $Dec<30$º that are not covered by deep multiband HST exposures. In particular, it will cover more than 1200 massive galaxies ($M_{B}<-14$ to select starforming hosts), where LRNe are most likely, and and whose YSGs comprise 96% of our sample. Assuming the same distance distribution of YSG candidates as the one in this study, this represents a 130% increase of our $M_{*}>6M_{\odot}$ sample. It is important to note that some progenitors detected only a few years before the LRN are likely already undergoing mass transfer and can be considered precursors (this is the case of e.g. AT2021blu, imaged both 15 years and 2 years prior to the transient, showing a 1-mag brightening between these two epochs Pastorello et al. 2023).
@@ 6 Conclusion
Our study identified 233,336 yellow supergiant candidates in 355 galaxies, a significant increase over previous research. The resulting sample of candidates was cross-matched with the TNS and the ongoing surveys ZTF, BlackGEM and MeerLICHT. Candidates exhibiting outstanding variability were identified and analysed. In particular, we identified 12 precursor candidates based on their consistent brightening over the past few years. Their spectroscopic followup and identification will be the subject of an upcoming work.
@@ Acknowledgements.
H. T. and N. B. acknowledge to be funded by the European Union (ERC, CET-3PO, 101042610). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. We acknowledge the extensive use of the MAST database to conduct this study. Based on observations with the MeerLICHT telescope. MeerLICHT is built and run by a consortium consisting of Radboud University, the University of Cape Town, the South African Astronomical Observatory, the University of Oxford, the University of Manchester and the University of Amsterdam. MeerLICHT is hosted by SAAO. Based on observations with the BlackGEM telescopes. BlackGEM is built and run by a consortium consisting of Radboud University, the Netherlands Research School for Astronomy (NOVA), and KU Leuven with additional support from Armagh Observatory and Planetarium, Durham University, Hamburg Observatory, Hebrew University, Las Cumbres Observatory, Tel Aviv University, Texas Tech University, Technical University of Denmark, University of California Davis, the University of Barcelona, the University of Manchester, University of Potsdam, the University of Valparaiso, the University of Warwick, and Weizmann Institute of science. BlackGEM is hosted and supported by ESO.

### v2

@@ 3.2 Contamination fraction
To evaluate the contamination fraction from intrinsic extinction and foreground sources, we considered various types of contaminants. Foreground contamination was quantified using the TRILEGAL star counts. According to the method detailed in Section 2.4, approximately 5% of the sources in our YSG sample are expected to be foreground contaminants (including sources having high Gaia proper motion). The foreground contamination mainly affects galaxies at low Galactic latitudes ($|b|<25$º) and 75% of the galaxies have less than 10% of their YSG candidates being foreground, as shown in Figure 10 (right panel). As illustrated in this Figure, galaxies closer to the Galactic plane or toward the Galactic center exhibit higher contamination fractions. After using Gaia proper motions to clean the sample, removing 6,623 sources (4.3% of the sample), we therefore estimate the remaining foreground contamination to be $\lesssim$ 1%.
@@
Finally, contaminants such as quasars, Cepheids and luminous blue variables were identified and removed using Gaia and Simbad databases. These known contaminants represent a minority of the sample across all distances (Figure 7). These sources are typically brighter in both observed and absolute magnitudes, as shown in Figure 12. The clean sample of YSG contains 146,502 sources.
@@ 4.1 Completeness for known LRN progenitors
Table 5: Retrieval status of known LRN progenitors in the YSG sample.
@@
 [^15] LRN progenitors identified in the literature, and their retrieval status in our YSG sample. References: [1] MacLeod et al. 2017, [2] Pastorello et al. 2021a, [3] Pastorello et al. 2021b, [4] Blagorodnova et al. 2017, [5] Blagorodnova et al. 2021, [6] Cai et al. 2022 [7] Pastorello et al. 2023, [8] Smith et al. 2016a.
Many extragalactic LRNe discovered during the last decade had pre-outburst images taken several years before the transient (e.g. MacLeod et al. 2022 and references therein). For eight of them, listed in Table 5, their host galaxy is in our sample, allowing us to assess the rate of retrieved progenitors in our sample of YSG candidates. Three out of eight progenitors are missing in our sample. For AT2015dl, the transient is located outside the footprint of the HST observations. For NGC4490-OT, although the progenitor was measured and studied by Smith et al. (2016b), its apparent magnitude of $m_{\mathrm{F606W}}=23.58\pm$0.24 was too faint to be retrieved as an HSCv3 source (the faintest retrieved sources in a 30 arcsec radius circle around the progenitor position were 23.2 mag). For AT2019zhd, the progenitor had a low luminosity ($M_{\mathrm{F555W}}=0.21\pm 0.14$, Pastorello et al. 2021a), placing it in the most populated region of the Hertzsprung gap. Consequently, it could not be selected after the Gaussian mixture cut. Moreover, its proximity to the edge of the field of view prevents this source from being retrieved as a MAST-catalogued source, and its magnitude error would either way exclude it from our sample. Considering these results, our sample appears to represent a relatively comprehensive selection of LRN progenitors, with a $\sim$70completeness level in regions covered by HST.
@@ 4.2 Progenitors of past and ongoing transients: crossmatch to TNS
M 31 21.4 2016fbx 0.27 Clear 17.7 -6.7 Nova
@@
 [^16] The last column shows the type according to TNS with eventual corrections (V*: variable star). Objects are sorted by TNS name.
Our YSG sample can be used to find possible progenitors for past transients and further analyse them. To this end, we crossmatched our YSG candidates to the TNS public objects as of 2024 August 31st, with a match radius of 0.6”. The resulting list of 24 transients is detailed in Table 6. About half of these transients are already classified and some are well-studied in the literature (e.g. SN2016bau: Aryan et al. 2021. SN2024ggi: Pessi et al. 2024; Jacobson-Galán et al. 2024). For those, our HST matches may provide useful pre-outburst luminosity levels. The TNS report of AT2019ejn mentions multiple discoveries of this transient, suggesting it may be an LBV. For three other transients (AT2020aaqy, AT2021tmm and AT2021ytf), the $F814W$ magnitude is brighter than the TNS discovery magnitude, making them likely to also be variable stars. For nine other transients, the type was not straightforwardly identified in TNS and they deserve a detailed analysis of their light curves and HST progenitors. Therefore, we inspected HST colour images for every candidate, using the Hubble Legacy Archive (HLA) website [^17] https://hla.stsci.edu/hlaview.html. Cutouts of 10 arcsec side-to-side are displayed in Figure 13 and the position of the transient is marked by a circle.
@@
Figure 14: ZTF forced photometry light curves of the 9 candidates resulting from our TNS cross-match. The blue areas indicate the period that was used to set the baseline flux at the location of these transients.
@@
AT2022llt: this transient matches a $>20M_{\odot}$ YSG with a separation of 0.06”. Although ZTF forced photometry does not reveal any long-term behavior, the ATLAS light curve shows a slow brightening of the source in recent years (Figure 16, top-middle panel). Image inspection shows that source confusion is unlikely for this source (also supported by its CI=1.05). Followup of this source is encouraged.
@@ 4.3 Other precursor candidates
Table 7: Precursor candidates identified in this study.
@@
 [^20] Precursor candidates (sorted by galaxy name) identified in ZTF, MeerLICHT, or BlackGEM. The last column gives the slope of the best-fit linear trend to the ATLAS light curve.
@@ 5.1 Comparison of our YSG candidates to spectroscopic samples
In addition to using proper motions, variability is an independent criterion that helps distinguishing a YSG from a foreground star. For example, massive YSGs have been known to vary erratically compared to low-mass dwarfs. The bottom panel of Figure 17 shows the distribution of the Gaia-measured dGmag, the difference between brightest and faintest $G$-band magnitude observed during the course of the survey. Interestingly, stars classified as foreground but with a low PM exhibit variability on average 6 times greater than those with proper motion data (average dGmag=0.2 instead of 0.033, Figure 17). This suggests a misclassification of these stars, which may be actual members of M31. To test this hypothesis, we model the distribution of dGmag as a sum of two distributions: one for high-PM sources identified as foreground, and one for low-PM sources identified as YSGs and RSGs. We can then estimate the contribution of YSGs and RSGs within low-PM foreground-classified stars. The fraction of these candidate members is 68 $\pm$ 11%. This suggests that the residual contamination within low-PM stars with $V<18.5$ is actually $\sim$ 8% and not 23%. Furthermore, if our sample were to be used to identify progenitors of variable or transient events, the expected value of dGmag would be relatively large, further reducing the probability of association with a foreground source. On the other hand, although we expect most background AGN to be excluded in our sample selection (Section 2.4), some unresolved AGN undetected in X-rays may still produce some alerts in future deep transient surveys.
@@
Finally, we can quantify the completeness of our YSG catalogue in different magnitude bins. Using for reference the samples of YSG candidates in M31 from Drout et al. (2009) and Gordon et al. (2016) (see the corresponding CMD in Figure 18), and applying the same proper motion cut as done in our study, we obtain the completeness estimates shown in Figure 19. While the brightest stars are not always detected in third-party HST catalogues (and surveys other than HST are well-adapted to recover those), the fraction of recovered YSGs among recovered HST sources consistently remains around 60%. The lost fraction is primarily due to sources with high CI or located in the dotted-line rectangle shown in Figure 18. The latter occurs because our conservative threshold, designed to exclude the main-sequence and some extinguished main-sequence stars, also omits YSGs in the bluer part of the CMD.
@@ 5.2 Impact of photometric uncertainties and source confusion
One method to quantify source confusion is by using the concentration index (CI), defined as the difference in magnitude between two apertures, typically a smaller aperture of 0.05” and a larger aperture of 0.15” (for ACS and WFC3/UVIS), normalised such that its distribution peaks at 1. Objects with CI values around 1.0 are likely to be stars, while those with significantly higher CI values are likely to be extended sources or the confusion of several stars. Figure 20 presents the distribution of YSG candidates in the CI – $m_{F814W}$ plane. At bright magnitudes, the CI is close to 1, showing that sources are point-like and source confusion is not significant. The faint sources are more numerous and thus denser on the sky, especially in distant galaxies. Therefore, they are more affected by source confusion and can have CI values close to 2. Such sources are currently excluded from our sample, which is restricted to CI values $<1.5$, thereby reducing its completeness. A future release of the catalog, incorporating DOLPHOT photometry (Dolphin, 2016), is anticipated to address this issue more effectively.
Similarly, photometric uncertainties can bias the selection of YSGs in HST observations. These uncertainties arise from various sources, such as photon noise, background subtraction errors, and instrumental effects. When dealing with YSG candidates, identified based on their position in the CMD, even small errors in photometry can shift stars into or out of the YSG region. For example, in the NGC 4455 galaxy, an edge-on spiral galaxy located 7.3 Mpc away, observations were conducted using the $F814W$ filter for 1030 seconds. These parameters are close to the median values of our sample of galaxies with YSG candidates. When the colour-magnitude diagram is regenerated using magnitude values and errors modelled as a normal distribution, 9% of YSG candidates shift out of the selection region. Conversely, a similar number of previously unselected stars are shifted into the YSG selection region. Likewise, the variability of YSG can cause them to move out of the selection region; however, this variability primarily occurs over short timescales and within specific regions of the HR diagram, such as the instability strip (e.g., Evans 1993).
@@ 5.3 Search for LRN Progenitors and Precursors with LSST
Conversely, the study projects that LSST, over its 10-year survey, could detect around 50 red supergiant progenitors and several yellow and blue supergiants. It estimates that progenitors of Type Ic supernovae would be detectable if they are brighter than -4.0 magnitudes in the LSST $i$ band, respectively. Given their similar spectral type (A5) compared to YSGs, we assume a similar performance for YSG detection: we therefore expect that 77% of our YSG sample will be redetected by LSST, for sources having $M_{F814W}<-4$ (or $M_{*}>6M_{\odot}$, Figure 9). This will provide exquisite variability constraints for sources of $r<23$, allowing us to identify new LRN precursors, given LSST repeated observations and the long timescales of these events. Such early precursor discoveries will be essential to schedule precursor spectroscopy, allowing us to probe the line profiles, outflow structure and density (and therefore mass loss), ionization structure, and possible shocks in the system (e.g. Pejcha et al. 2016; Molnar et al. 2017; Blagorodnova et al. 2021).
Furthermore, LSST can expand the progenitor sample for the 1235 galaxies within 20 Mpc and at $Dec<30$º that are not covered by deep multiband HST exposures. In particular, it will cover more than 1200 massive galaxies ($M_{B}<-14$ to select starforming hosts), where LRNe are most likely [^21] Every extragalactic LRN discovered in the last decade belonged to a massive starforming host, the least massive host being the Spider galaxy with $M_{B}=-16.7$ (Pastorello et al., 2023)., and and whose YSGs comprise 96% of our sample. Assuming the same distance distribution of YSG candidates as the one in this study, this represents a 130% increase of our $M_{*}>6M_{\odot}$ sample. It is important to note that some progenitors detected only a few years before the LRN are likely already undergoing mass transfer and can be considered precursors (this is the case of e.g. AT2021blu, imaged both 15 years and 2 years prior to the transient, showing a 1-mag brightening between these two epochs Pastorello et al. 2023). Ongoing large-scale, deep surveys such as Euclid (Laureijs et al., 2011) are also of significant interest for mapping stellar populations and identifying potential transient progenitors in nearby galaxies (Bonanos et al., 2024).
@@ 6 Conclusion
Our study identified 146,502 yellow supergiant candidates in 353 galaxies, a significant increase over previous research. The resulting sample of candidates was cross-matched with the TNS and the ongoing surveys ZTF, BlackGEM and MeerLICHT. Candidates exhibiting outstanding variability were identified and analysed. In particular, we identified 12 precursor candidates based on their consistent brightening over the past few years. Their spectroscopic followup and identification will be the subject of an upcoming work.
@@ Acknowledgements.
H. T. and N. B. acknowledge to be funded by the European Union (ERC, CET-3PO, 101042610). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. PJG is supported by NRF SARChI grant 111692. We thank Zeljko Ivezic and the anonymous referee for their valuable comments, which significantly enhanced the quality of this paper. We thank Rick White and Bernie Shiao for their assistance in retrieving the HSCv3 data. We acknowledge the extensive use of the MAST database to conduct this study. Based on observations with the MeerLICHT telescope. MeerLICHT is built and run by a consortium consisting of Radboud University, the University of Cape Town, the South African Astronomical Observatory, the University of Oxford, the University of Manchester and the University of Amsterdam. MeerLICHT is hosted by SAAO. Based on observations with the BlackGEM telescopes. BlackGEM is built and run by a consortium consisting of Radboud University, the Netherlands Research School for Astronomy (NOVA), and KU Leuven with additional support from Armagh Observatory and Planetarium, Durham University, Hamburg Observatory, Hebrew University, Las Cumbres Observatory, Tel Aviv University, Texas Tech University, Technical University of Denmark, University of California Davis, the University of Barcelona, the University of Manchester, University of Potsdam, the University of Valparaiso, the University of Warwick, and Weizmann Institute of science. BlackGEM is hosted and supported by ESO.